# ReQISC: A Reconfigurable Quantum Computer Microarchitecture and Compiler Co-Design

### v1

@@ 6.1.4. Hardware assumptions
The ReQISC microarchitecture is evaluated under representative coupling Hamiltonians, including $\mathrm{XY}$, $\mathrm{XX}$, and random couplings. Since $\mathrm{XY}$ coupling aligns with the mainstream flux-tunable superconducting platforms (e.g., Google’s Sycamore (arute2019quantum)), it is the default assumption for pulse-level benchmarking and fidelity experiments. Baselines are compared using their conventional, optimized pulse implementations for $\mathrm{CNOT}$, $\mathrm{iSWAP}$, and $\mathrm{SQiSW}$ gates under $\mathrm{XY}$ coupling (krantz2019quantum; arute2019quantum; huang2023quantum).
@@ 6.2. Performance of ReQISC microarchitecture
We quantitatively benchmark the ReQISC microarchitecture by computing the average pulse duration to synthesize $10^{5}$ Haar-random $\mathrm{SU}(4)$ unitaries (Table 2). For mainstream $\mathrm{XY}$-coupled platforms, ReQISC slashes the average synthesis cost to just $1.341\,g^{-1}$—a nearly five-fold improvement over the $6.664\,g^{-1}$ ($3\times\pi/\sqrt{2}\approx 6.664$ (krantz2019quantum)) required by conventional $\mathrm{CNOT}$ gate scheme. This superior efficiency is not limited to a specific hardware interaction; ReQISC also achieves an average duration of $1.178\,g^{-1}$ for XX coupling and $1.321\,g^{-1}$ for arbitrary random couplings. These results demonstrate that a full $\mathrm{SU}(4)$ ISA, when powered by our microarchitecture, is fundamentally more performant than conventional approaches. Furthermore, our scheme is so efficient that even when used to realize a fixed gate like $\mathrm{SQiSW}$, its performance is nearly on par (within a 1.61.6x factor) with that of the fully arbitrary $\mathrm{SU}(4)$ implementation.
@@ 6.3. Logical-level compilation
As detailed in Table 1, ReQISC delivers substantial improvements over all baselines across every key metric. Particularly, ReQISC consistently slashes the total pulse-level circuit duration by 4040%–9090% across all benchmarks, with an overall average of $68\%$ ($73\%$) for ReQISC-Red (ReQISC-Full). We note that the difference between the results for ReQISC-Red versus ReQISC-Full is much more pronounced for non-variational quantum programs, suggesting that they benefit more from local optimization via hierarchical synthesis. That aligns with our expectation since these arithmetic-logic-like programs exhibit versatile subcircuit patterns.
@@
Figure 10. topology-aware benchmarking for 1D-chain (upper) and 2D-grid (bottom) mapping. _logic: compilation on logical circuits without mapping; phys_opt: optimization on physical circuits by means of Qiskit’s topology-preserving O3 transpile. Green dots and blue triangles indicate the multiples of #2Q after optimized qubit mapping (hatched-green or hatched-blue) relative to logical circuits (light-green or light-blue). Dashed horizontal lines refer to the geometric-mean overheads.
@@
Figure 11. Trade-off between calibration overhead and #2Q reduction (ReQISC-Red v.s. ReQISC-Full). Each connected pair of points corresponds to a quantum program instance. Scaling trends remain consistent for larger circuits (not shown due to fewer data points).
@@
6.4. Topology-aware compilation
Figure 10 illustrates ReQISC’s performance with qubit mapping/routing, across a set of medium-scale benchmarks from each category. Results show that mirroring-SABRE outperforms conventional SABRE paired with ReQISC, with #2Q reduction up to 28.228.2% (23.723.7%) and on average 11.011.0% (15.715.7%) for 1D-chain (2D-grid) qubit mapping. This effect is also program-dependent—for instance, mirroring-SABRE incurs no additional #2Q overhead on qft_8 and tof_10. The advantage is even more pronounced when comparing the end-to-end ReQISC compilation against the SOTA $\mathrm{CNOT}$-based baseline (hatched-blue v.s. hatched-green in Figure 10). Specifically, the geometric mean of $\mathrm{SU}(4)$ overhead is 1.361.36x (1.091.09x) for 1D-chain (2D-grid) mapping; for $\mathrm{CNOT}$, it is 2.452.45x (1.791.79x). These results highlight the fundamental superiority of ReQISC for compiling to realistic hardware.
6.5. Calibration efficiency v.s. #2Q reduction
We evaluate the trade-off between calibration overhead and 2Q gate count reduction by comparing the number of distinct $\mathrm{SU}(4)$ gates in circuits compiled by ReQISC-Red and ReQISC-Full for benchmarks with #2Q up to $5000$. As shown in Figure 11, ReQISC-Red has negligible calibration overhead as expected, with distinct $\mathrm{SU}(4)$s below $10$. In field tests, only $56$ distinct 3Q subcircuit structures were found across all benchmarks except variational programs. For ReQISC-Full, calibration overhead is reasonable, typically below $200$. This demonstrates a practical and controllable trade-off for users seeking the highest performance, allowing them to balance gate fidelity with experimental complexity.
6.6. Ablation study and Breakdown analysis
@@ 6.6.1. Efficacy of specialized passes
To isolate the benefits of our co-designed compilation strategy, we perform ablation study for ReQISC-Full against Qiskit-$\mathrm{SU}(4)$, TKet-$\mathrm{SU}(4)$, and BQSKit-$\mathrm{SU}(4)$. As shown in Figure 12, ReQISC-Full still significantly outperforms these variants, confirming that its global program-aware and hierarchical synthesis passes are critical for effective $\mathrm{SU}(4)$ compilation. Notably, while BQSKit-$\mathrm{SU}(4)$ achieves a reasonable gate count reduction, it is impractical due to an explosion of distinct $\mathrm{SU}(4)$ gates, which implies an uncontrollable calibration overhead. In contrast, ReQISC-Full maintains a manageable, sub-linearly scaling set of distinct gates—an order of magnitude fewer in our tests (Figure 11).
@@ 6.6.2. Impact of DAG Compacting
Our breakdown analysis confirms the importance of the DAG compacting pass. We first observed that different circuit partitioning strategies had a negligible impact on performance (less than $2\%$ variation). The DAG compacting pass, however, is crucial. Comparing ReQISC-Full against a version without this pass (“ReQISC-NC”) reveals substantial gains. As shown in Figure 12, this single pass improves the #2Q reduction rate by up to $33\%$ (e.g., for rip_add benchmarks), demonstrating its effectiveness in exploiting approximate commutation rules.
Figure 13. Program fidelity $F$ and pulse duration $T$ comparison.
Figure 14. Compilation errors (a) and latency (b) comparison.
@@ 6.7. Fidelity experiment
We performed noisy simulations by QSim (quantum_ai_team_and_collaborators_2020_4023103) on twelve representative benchmarks, comparing ReQISC against a SOTA $\mathrm{CNOT}$-based workflow (TKet with SABRE mapping). Our model assumed a $\mathrm{XY}$-coupled device with the depolarizing channel appended to 2Q gates, where error rates were scaled proportionally to the gate duration ($\tau$) of each $\mathrm{CNOT}$ ($\tau_{0}=\pi/\sqrt{2}\,g^{-1};\,p_{0}=0.001$) or $\mathrm{SU}(4)$ operation ($\tau;\,p=p_{0}\tau/\tau_{0}$). As shown in Figure 13, ReQISC delivers substantially higher fidelity and faster execution across the board. At the logical level (all-to-all topology), it achieves an average 2.362.36x error reduction and a 3.063.06x execution speedup. These advantages become even more pronounced when mapping to topology-constrained devices, with error reductions of 3.183.18x (2D grid) and 3.343.34x (1D chain), and corresponding speedups of 4.304.30x and 4.554.55x, respectively.
@@ 6.8. Reliability and scalability
For benchmarks with up to eleven qubits, we verify the compiled circuits generated by baselines and the ReQISC compiler, by computing the compilation error quantified by circuit infidelity. Figure 14 (a) shows all compilers exhibit a similar level of compilation errors. This is expected, as the approximate synthesis is de facto “numerically exact” for every subcircuit synthesis, whose correctness is bounded by the IEEE754 machine precision $\approx 2.2\times 10^{-16}$. The extensive peephole optimizations (e.g., block consolidation, KAK decomposition) within Qiskit/TKet can only be at the level of machine precision as well. Note that both $x$ and $y$ axes use logarithmic scaling—for large circuits, ReQISC/Qiskit/TKet’s compilation errors are bounded much better than BQSKit.
The ReQISC compiler’s computational complexity is polynomial in program size, as its key passes have linear or polynomial complexity. Figure 14 (b) confirms that both ReQISC-Red and ReQISC-Full scale well in their runtime, despite being implemented in Python. ReQISC-Red is consistently more efficient than TKet (C++ backend) and BQSKit (Rust backend). ReQISC-Full, which performs heavy optimizations, offers performance competitive with BQSKit, demonstrating that its substantial gate count reduction is achieved with a practical and acceptable increase in compilation time.
@@ 7. Conclusion
$\mathrm{CNOT}$-based ISAs are the predominant choice for most hardware platforms. Although some continuous gate families have been proposed and experimentally demonstrated, they are generally impractical due to suboptimal or non-native implementation, excessive calibration overhead, and limited compilation strategies. The presented ReQISC uses the most expressive $\mathrm{SU}(4)$ ISA that is optimally realized by means of our microarchitecture. Our approach for the first time enables the straightforward implementation of any equivalent 2Q gate under arbitrary coupling Hamiltonians, with simple pulse control and minimal gate time. Supported by the end-to-end ReQISC compiler, we can strike an excellent balance between performance and calibration overhead, thereby challenging traditional notions of the two being fundamentally in contradiction.
@@
The implications of this work are significant for quantum computing both in the near and long term. For the fault-tolerant regime (preskill2025beyond; acharya2024quantum), our approach offers multifaceted advantages: it provides a pathway to faster and higher-fidelity logical qubits (e.g., via a $\sqrt{2}$x faster $\mathrm{CNOT}$ on XY-coupled hardware) and enables native support for Clifford operations like $\mathrm{iSWAP}$ and $\mathrm{SWAP}$ that proves critical for state-of-the-art QEC schemes (mcewen2023relaxing; eickbusch2024demonstrating; zhou2024halma; wu2022synthesis). By providing the architecture and compiler to leverage a richer, more expressive native gate set, ReQISC opens new avenues for designing performant and scalable quantum computers.

### v2

@@ 6.1.4. Hardware assumptions
The ReQISC microarchitecture is evaluated under representative coupling Hamiltonians, including $\mathrm{XY}$, $\mathrm{XX}$, and random couplings. Since $\mathrm{XY}$ coupling aligns with the mainstream flux-tunable superconducting platforms (e.g., Google’s Sycamore (Arute et al., 2019)), it is the default assumption for pulse-level benchmarking and fidelity experiments. Baselines are compared using their conventional, optimized pulse implementations for $\mathrm{CNOT}$, $\mathrm{iSWAP}$, and $\mathrm{SQiSW}$ gates under $\mathrm{XY}$ coupling (Krantz et al., 2019; Arute et al., 2019; Huang et al., 2023).
@@ 6.2. Performance of ReQISC microarchitecture
We benchmark the ReQISC microarchitecture by computing the average pulse duration to synthesize $10^{5}$ Haar-random $\mathrm{SU}(4)$s (Table 3). For mainstream $\mathrm{XY}$-coupled platforms, ReQISC slashes the average synthesis cost to just $1.341\,g^{-1}$—a nearly five-fold improvement over $6.664\,g^{-1}$ ($3\times\pi/\sqrt{2}\approx 6.664$ (Krantz et al., 2019)) required by conventional $\mathrm{CNOT}$ realization. This efficacy is consistent across hardware interactions, with average durations of $1.178\,g^{-1}$ for $\mathrm{XX}$ coupling and $1.321\,g^{-1}$ for arbitrary random couplings. These results demonstrate that a full $\mathrm{SU}(4)$ ISA, when powered by our microarchitecture, fundamentally outperforms conventional approaches. Our scheme is so efficient that even when used to realize a fixed gate like $\mathrm{SQiSW}$, its performance is nearly on par (within a factor of $1.6$) with that of the fully arbitrary $\mathrm{SU}(4)$.
@@ 6.3. Logical-level compilation
As detailed in Table 2, ReQISC delivers substantial improvements over all baselines across every key metric. Particularly, ReQISC consistently slashes the total pulse-level circuit duration by 4040%–9090% across all benchmarks, with an overall average of $68\%$ ($71\%$) for ReQISC-Eff (ReQISC-Full). The difference between the results for ReQISC-Eff versus ReQISC-Full is much more pronounced for non-variational, arithmetic-logic programs, suggesting that they benefit more from local optimization via hierarchical synthesis. That aligns with our expectation since they exhibit versatile subcircuit patterns.
@@
Figure 12. topology-aware benchmarking for 1D-chain (upper) and 2D-grid (bottom) mapping. xxx_logic: logical circuits (pre-mapping); phys_opt: physical circuits with Qiskit O3 topology-preserving optimization. Green dots and blue triangles indicate the multiples of #2Q after optimized qubit mapping (hatched-green or hatched-blue) relative to logical circuits (light-green or light-blue). Dashed horizontal lines refer to the geometric-mean overheads.
@@
Figure 13. Calibration efficiency and trade-off with #2Q reduction (ReQISC-Eff vs. ReQISC-Full). (a) Each connected pair of points corresponds to a benchmark program. Scaling trends remain consistent for larger circuits (not shown due to fewer data points). (b) Distinct $\mathrm{SU}(4)$ count distributions.
6.4. Topology-aware compilation
Figure 12 illustrates ReQISC’s performance with qubit mapping/routing, across medium-scale benchmarks per category. Results show that mirroring-SABRE outperforms conventional SABRE paired with ReQISC, with #2Q reduction up to 28.228.2% (23.723.7%) and on average 11.011.0% (15.715.7%) for 1D-chain (2D-grid) mapping. This effect is also program-dependent—for instance, mirroring-SABRE incurs no additional #2Q overhead on qft_8 and tof_10. The advantage is even more pronounced when comparing the end-to-end ReQISC compilation against the SOTA $\mathrm{CNOT}$-based baseline (hatched-blue vs. hatched-green). Specifically, the geometric mean of $\mathrm{SU}(4)$ overhead is 1.361.36x (1.091.09x) for 1D-chain (2D-grid) mapping; for $\mathrm{CNOT}$, it is 2.452.45x (1.791.79x). These results underscore the ReQISC’s superiority in topology-aware compilation.
6.5. Calibration efficiency and tradeoff discussion
We evaluate the trade-off between calibration overhead and 2Q gate count reduction by comparing the number of distinct $\mathrm{SU}(4)$ gates in circuits compiled by ReQISC-Eff and ReQISC-Full for benchmarks with #2Q up to $5000$. As shown in Figure 13, ReQISC-Eff has negligible calibration overhead as expected, with less than 10 distinct $\mathrm{SU}(4)$s. For ReQISC-Full, the number of distinct SU(4)s is below $200$ (Figure 13(a)), while more than three quarters of ReQISC-Full-compiled programs exhibits less than $20$ distinct 2Q gates (Figure 13(b)). This demonstrates a practical and controllable trade-off for users seeking the highest performance, allowing for a balance between gate fidelity with experimental complexity.
6.5.1. System-level calibration scalability
A critical concern for adopting continuous ISAs is whether the calibration cost scales prohibitively with the number of user programs. We argue that ReQISC avoids this potential “calibration explosion” through several synergistic mechanisms.
First, the structural regularity of quantum algorithms ensures a “bounded template library”. Implementing arbitrary 2Q gates does not require supporting every possible unitary in practice; instead, most programs require only a few distinct $\mathrm{SU}(4)$ operations. This implies that a system provider only needs to maintain a calibrated library of a few dozen “standard gates” to serve a vast range of applications, keeping the marginal cost of supporting new programs negligible.
Second, our framework utilizes “model-based parameter generation”. Since the ReQISC microarchitecture is physically derived from the coupling Hamiltonian, once the underlying hardware model (e.g., $g$ and drive strengths) is characterized—a standard periodic maintenance task—control parameters for any new $\mathrm{SU}(4)$ can be analytically generated or interpolated with high initial fidelity (Chen et al., 2024, 2025). Recent experimental work by Kakkar et al. (2025) further validates this approach, demonstrating that a continuous gate family can be implemented with minimal to zero extra calibration cost by characterizing only a small subset of the parameter space.
Furthermore, we emphasize that ReQISC’s “reconfigurability” aligns with existing experimental standards. High-fidelity computation typically necessitates recalibration before running any nontrivial program, even on fixed-ISA systems. The perception of high cost for rich ISAs is largely a byproduct of them not yet being mainstream. In practice, the total calibration cost scales linearly with distinct $\mathrm{SU}(4)$ count. It is experimentally feasible, as evidenced by the successful calibration of six distinct gates in Chen et al. (2025).
While the number of distinct gates affects calibration overhead, a larger gate set may also introduce more complex parallel crosstalk calibration. Fortunately, our unified control framework for all $\mathrm{SU}(4)$ operations significantly simplifies this process. This is demonstrated by the successful calibration of six distinct gates in Chen et al. (2025), as previously noted, and could be further streamlined by the techniques proposed in Kakkar et al. (2025).
6.6. Ablation study and Breakdown analysis
@@ 6.6.1. Efficacy of specialized passes
To isolate the benefits of our co-designed compilation strategy, we conduct the ablation study for ReQISC-Full against Qiskit-$\mathrm{SU}(4)$, TKet-$\mathrm{SU}(4)$, and BQSKit-$\mathrm{SU}(4)$. As shown in Figure 14, ReQISC-Full still significantly outperforms these variants, confirming that its global program-aware and hierarchical synthesis passes are critical for effective $\mathrm{SU}(4)$ compilation. Notably, while BQSKit-$\mathrm{SU}(4)$ achieves a reasonable gate count reduction, it is impractical due to an explosion of distinct $\mathrm{SU}(4)$ gates, which implies an uncontrollable calibration overhead. In contrast, ReQISC-Full maintains a manageable, sub-linearly scaling set of distinct gates—an order of magnitude fewer in our tests (Figure 13).
@@ 6.6.2. Impact of DAG Compacting
Our breakdown analysis confirms the importance of the DAG compacting pass. We first observed that different circuit partitioning strategies had a negligible impact on performance (less than $2\%$ variation). The DAG compacting pass, however, is crucial. Comparing ReQISC-Full against a version without this pass (“ReQISC-NC”) (“NC” denotes no-compacting) reveals substantial gains. As shown in Figure 14, this single pass improves the #2Q reduction rate by up to $33\%$ (e.g., for rip_add benchmarks), demonstrating its effectiveness in exploiting approximate commutation rules.
@@ 6.7. Fidelity experiment
Figure 15. Program fidelity $F$ and pulse duration $T$ comparison through noisy simulation.
We performed noisy simulations using QSim (team and collaborators, 2020) on twelve representative benchmarks, comparing ReQISC against a SOTA $\mathrm{CNOT}$-based workflow (TKet with SABRE mapping). Our model assumed a $\mathrm{XY}$-coupled device with the depolarizing channel appended to 2Q gates, where error rates were scaled proportionally to the gate duration ($\tau$) of each $\mathrm{CNOT}$ ($\tau_{0}=\pi/\sqrt{2}\,g^{-1};\,p_{0}=0.001$) or $\mathrm{SU}(4)$ operation ($\tau;\,p=p_{0}\tau/\tau_{0}$). As shown in Figure 15, ReQISC delivers substantially higher fidelity and faster execution across the board. At the logical level (all-to-all topology), it achieves an average 2.362.36x error reduction and a 3.063.06x execution speedup. These advantages become even more pronounced when mapping to topology-constrained devices, with error reductions of 3.183.18x (2D grid) and 3.343.34x (1D chain), and corresponding speedups of 4.304.30x and 4.554.55x, respectively.
@@ 6.8. Reliability and scalability
Figure 16. Compilation errors (a) and latency (b).
For benchmarks comprising up to eleven qubits, we verify the circuits generated by baselines and the ReQISC compiler by computing the compilation error quantified by circuit infidelity. Figure 16(a) shows that all compilers exhibit a similar level of compilation errors. This is expected, as the approximate synthesis is de facto “numerically exact” for every subcircuit synthesis, whose correctness is bounded by the IEEE754 machine precision $\approx 2.2\times 10^{-16}$. The extensive peephole optimizations (e.g., block consolidation, KAK decomposition) within Qiskit/TKet can only be at the level of machine precision as well. Note that both $x$ and $y$ axes use logarithmic scaling—for large circuits, ReQISC/Qiskit/TKet’s compilation errors are bounded much better than BQSKit.
The ReQISC compiler’s computational complexity is polynomial in program size, as its key passes have linear or polynomial complexity. Figure 16(b) confirms that both ReQISC-Eff and ReQISC-Full scale well in their runtime, despite being implemented in Python. ReQISC-Eff is consistently more efficient than TKet (C++ backend) and BQSKit (Rust backend). ReQISC-Full, which performs heavy optimizations, offers performance competitive with BQSKit, demonstrating that its substantial gate count reduction is achieved with a practical and acceptable increase in compilation time.
@@ 7. Conclusion
Traditional quantum hardware relies predominantly on $\mathrm{CNOT}$-based ISAs. Although some continuous gate families have been proposed and experimentally demonstrated, they are generally impractical due to suboptimal or non-native implementation, excessive calibration overhead, and limited compilation strategies. ReQISC addresses these limitations with a hardware-native $\mathrm{SU}(4)$ ISA, optimally realized by the underlying microarchitecture across various hardware. For the first time, our approach enables the straightforward implementation of any equivalent 2Q gate under arbitrary coupling Hamiltonians, with simple pulse control and minimal gate time. Supported by the end-to-end ReQISC compiler, we can strike an excellent balance between performance and calibration overhead, thereby challenging the traditional notion that performance and calibration overhead are fundamentally in contradiction.
@@
The implications of this work are significant for quantum computing both in the near and long term. For the fault-tolerant regime (Preskill, 2025; Acharya et al., 2024), our approach offers multifaceted advantages: it provides a pathway to faster and higher-fidelity logical qubits (e.g., via a $\sqrt{2}$x faster $\mathrm{CNOT}$ on XY-coupled hardware) and enables native support for Clifford operations like $\mathrm{iSWAP}$ and $\mathrm{SWAP}$ that proves critical for SOTA QEC schemes (McEwen et al., 2023; Eickbusch et al., 2025; Zhou et al., 2024; Wu et al., 2022). By providing the architecture and compiler to leverage a richer, more expressive native gate set, ReQISC opens new avenues for designing performant and scalable quantum computers.
Acknowledgements.
We thank all anonymous reviewers for their thoughtful comments. This research was partially conducted by AI Chip Center for Emerging Smart Systems (ACCESS), supported by the InnoHK initiative of the Innovation and Technology Commission of the Hong Kong Special Administrative Region Government. It was also supported partially by Research Grants Council of Hong Kong SAR (#16213824 & #16212825). Portions of this work were conducted while Z. Y. was attending UC Santa Barbara and visiting Tsinghua University, and while D. D. was at Tsinghua University. D. D. would like to thank God for all of His provisions.
Artifact Appendix
Abstract
ReQISC couples the Regulus SU(4)-native compiler with the genAshN microarchitecture (time-optimal SU(4) gate scheme). The artifact bundles: (1) the compiler implementation in regulus/; (2) the genAshN gate-duration model in microarch/genAshN; (3) the benchmark suite in benchmarks/; and (4) automation in artifact/ (Makefile + scripts + notebooks) to regenerate the paper’s tables and figures. The Makefile executes the full compilation pipeline, aggregates metrics, and emits LaTeX tables and PDF figures from the produced CSVs.
Artifact check-list (meta-information)
  • Algorithm: SU(4)-native template-based synthesis, hierarchical approximate synthesis, mirroring-SABRE qubit mapping; genAshN gate scheme solving
  • Program: Python 3.8+ (Regulus compiler and scripts)
  • Transformations: template-based synthesis, hierarchical SU(4) synthesis, mirroring, decomposition, circuit concatenation, circuit-DAG conversion
  • Data set: benchmarks/Type-I, benchmarks/Type-II QASM/JSON quantum programs with high-level subroutines
  • Run-time environment: Linux/macOS, Python virtualenv
  • Hardware: x86_64 with >>16 GB RAM recommended; CPU-bound
  • Run-time state: deterministic; no RNG required
  • Execution: Makefile targets wrapping Python scripts
  • Metrics: two-qubit gate count, two-qubit depth, pulse duration
  • Output: CSVs in artifact/results, QASM in artifact/output, LaTeX in artifact/notebooks/eval_compiler_main
  • Experiments: compiler benchmarking vs. Qiskit, TKET, BQSKit; pulse-duration comparison
  • Prep time: <<10 minutes for environment setup
  • Experiment time: Demo <1<1 minutes; Full make results several hours (Regulus-Eff/Full + Baselines (Qiskit/TKet/BQSKit))
  • Publicly available?: Yes
  • Code licenses?: Apache 2.0
  • Data licenses?: Same as code (benchmarks included)
  • Workflow automation?: GNU Make
  • Archived?: https://zenodo.org/records/18163249 with DOI 10.5281/zenodo.18163249
Description
How to access
Artifact is accessible via the Zenodo archive https://zenodo.org/records/18163249.
Main entry points:
  • regulus/: Regulus compiler implementation.
  • microarch/genAshN/: Microarchitecture implementation (gate scheme)
  • benchmarks/: Benchmark circuits with metadata.
  • artifact/scripts/: Command-line interface scripts.
  • artifact/notebooks/: Jupyter notebooks for figures and tables.
  • artifact/Makefile: Automation of experiment workflow.
Hardware dependencies
Any modern x86_64 CPU or Apple Silicon; 8 cores and >>16 GB RAM recommended. No GPU/accelerator required.
Software dependencies
Python 3.8+, pip, virtualenv. Packages listed in requirements.txt (qiskit, bqskit, tket, rustworkx, tqdm, etc.). Dockerfile provided for containerized runs.
Data sets
Benchmarks, with metadata, data preparation and pre-processing scripts in benchmarks/. No external downloads.
Models
The ReQISC microarchitecture design protocols (Algorithm 1) implemented in microarch/genAshN provides duration computation model for pulse-level comparisons; no training required.
Installation
  1. (1)
    Create env:
        cd artifact
        make env
        source .venv/bin/activate
  2. (2)
    (Optional) Docker: docker build -t regulus . then run a container and execute the same Make targets inside.
Experiment workflow
  1. (1)
    Sanity run (minutes): make demo (Regulus reduced + Qiskit on category alu).
  2. (2)
    Full regeneration (hours): make results. Runs Regulus (reduced and full), Qiskit, TKET, and BQSKit on all categories; writes
    • QASM to artifact/output/<compiler>/<category>/
    • result to artifact/results/<compiler>/<category>/
    • aggregated to artifact/results/result_<compiler>.csv
  3. (3)
    Tables only: make tables regenerates from the stored CSVs to compilation_result.tex.
  4. (4)
    Figures: Run the scripts in artifact/notebooks/ (each folder maps to a figure suite).
Evaluation and expected results
  • Circuit-level metrics: Two-qubit gate/depth reductions, contained in artifact/results/result_*.csv contain two-qubit gate/depth reductions matching the paper. Success: regenerated, non-empty CSVs.
  • Pulse metrics: pulse_comparison.py under artifact/ summarizes duration reductions; pulse_comparison.tex formats the table.
  • LaTeX tables: compilation_result.tex matches manuscript tables.
Experiment customization
  • Limit categories: edit CATEGORIES in artifact/Makefile or invoke scripts/bench.py on a specific directory.
  • Runtime vs. quality: set TRIALS=<int> when invoking make to control Regulus hierarchical synthesis trials.
  • Topology-aware runs: pass --device chain or --device grid to bench.py if desired.
Notes
  • If TKET or BQSKit are unavailable on your platform, regenerate tables from provided CSVs via make tables.
  • Delete artifact/output/ between runs by executing make clean-output to avoid mixing fresh and archived QASM.